OpenStudy (anonymous):

The point (3, -4) is on the terminal side of an angle . What is cos theta ?

OpenStudy (anonymous):

*Is on the terminal side of of angle theta

OpenStudy (anonymous):

the answers choices: -3/5 3/5 -3/4 3/4 3

OpenStudy (anonymous):

This is the right triangle. \[\cos \theta=\frac{ adjacent }{ hypotenuse }\] The adjacent side is 3. Find the hypotenuse 

OpenStudy (anonymous):

how? @peachpi

OpenStudy (anonymous):

is it 25?

OpenStudy (anonymous):

the pythagorean theorem a² + b²=c²

OpenStudy (anonymous):

you have to take the square root of 25

OpenStudy (anonymous):

so it is 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would the answer to my question be 3/5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous):

you're welcome